# On continuity of state-dependent utilities

### v1

@@ 3 Statement of the main results
    a probability measure $\mathbb{P}$ on ${\mathcal{F}}$,
@@
    a ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ with $\mathbb{u}(\omega,\cdot)$ being a continuous and strictly increasing function on $\mathbb{R}$ for every $\omega\in\Omega$ and $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$,
@@
In our second result we apply Theorem 3.1 to obtain an explicit formula for conditional Chisini means. Given a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\rightarrow\mathbb{R}$ and a random variable $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, DM23 proposes sufficient conditions to guarantee the existence of a solution $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ to the infinite dimensional system of equations $T(f\mathbf{1}_{A})=T(g\mathbf{1}_{A})$ for any $A\in{\mathcal{G}}$. The family of solutions is the so-called conditional Chisini mean, denoted by $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$. If compared with (1), it evidently appears that $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is the nonlinear counterpart of the standard conditional expectation. It is known that the problem of existence of nonlinear conditional expectation has been addressed also in CoquetPeng02 , yet the provided solution relies on a Brownian framework which in turn demands for more technical assumptions.
Embedding such a problem in a decision theoretical setup, we could rephrase the issue as follows: given a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ then the certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is $a\in\mathbb{R}$ such that $a\sim f$. Similarly if the agent is provided with additional information, represented by the $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, then the conditional certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is any $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that
$f\mathbf{1}_{A}\sim g\mathbf{1}_{A}\quad\forall\;A\in{\mathcal{G}}.$ (4)
Loosely speaking, $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ can be considered as a ${\mathcal{G}}$-measurable counterpart of the ${\mathcal{F}}$-measurable function $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ (see also FM11cce and MM21 for applications of the conditional certainty equivalent).
Since this paper deals with state dependent preferences we shall directly refer to the conditional Chisini mean as the class
[…]
Theorem 3.2
[…]
  • 1.
    a function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$, with
@@
      $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ continuous and strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$,
@@
and recall that, under the assumptions of Theorem 3.2, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is unique up to $\succeq$-null events (see DM23 , Theorem 3.2). To clarify the statement of Theorem 3.2, it seems important at this stage to point out that the use of $\Phi_{\mathcal{G}}$ in (6) conceals some technical subtleties. To begin with, the dependence of the image of $\mathbb{u}_{\mathcal{G}}$ on $\omega$ requires cautious handling of its domain. Additionally, for $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, the nature of $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ implies that it may not be an element of $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ itself. While the case of a finitely generated $\sigma$-algebra is rather simple (as we shall depict in Example 3.3 below), we shall dedicate a comprehensive treatment to the issue for an arbitrary $\sigma$-algebra ${\mathcal{G}}$ in the proof postponed in Section 5.
Example 3.3
@@
For any fixed $A_{i}$ we notice that the image of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$, namely $R_{i}:=\mathrm{Im}(\mathbb{u}_{\mathcal{G}}(\omega,\cdot))$ does not depend on $\omega\in A_{i}$. Given the generalized inverse $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ defined in Theorem 3.2, it is crucial to observe that the map $x\mapsto\Phi_{\mathcal{G}}(\omega,x)=\mathbb{u}_{\mathcal{G}}^{-1}(\omega,x)$ is well defined as a standard inverse function for $\omega\in A_{i}$ and $x\in R_{i}$ (as a result of strict monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$). In particular if we now choose any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and set $h:=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ then $h(\omega)\in R_{i}$ for $\omega\in A_{i}$ and indeed
@@
on the set $A_{i}$. This can be readily verified following the procedure in Section 5, step 4.
4 Proof of Theorem 3.1

### v2

@@ 3 Solution to an open question in Wakker and Zank (1999)
    a probability measure $\mathbb{P}$ on $(\Omega,{\mathcal{F}})$ such that ${\mathcal{N}}_{\succeq}=\left\{A\in{\mathcal{F}}\middle|\mathbb{P}(A)=0\right\}$,
@@
    a function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ with $\mathbb{u}(\omega,\cdot)$ strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$ and $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$,
@@
The previous theorem leaves an interesting issue unsolved as declared by the authors [24], page 29:
“[…] Pointwise continuity is not overly restrictive because continuity of each state-dependent utility implies pointwise continuity. Whether the reversed implication holds, i.e., whether continuity of the functions $\mathbb{u}$, in Statement (i) of Theorem [^6] Referred to the original statement in [24] 12 holds (after appropriate modification on null events), is an open question to us.”
To the best of our knowledge this question did not find an answer in the subsequent literature. Among the principal aims of the present paper, Theorem 7 stated below provides the construction of a pointwise continuous version of the state-dependent utility $\mathbb{u}$ [^7] Notice that in Appendix A of [17] a somewhat cumbersome notion of continuity was introduced to overcome some technical hurdles in the proof of the main result. Guaranteeing the existence of a continuous $\mathbb{u}$ allows to easily circumvent issues of this kind..
Theorem 7.
Assume that ${\mathcal{F}}$ contains at least three disjoint non-null events and $\succeq$ satisfies (SM), (PC) and (ST). Then $\mathbb{u}$ in Theorem 6 can be chosen to be ${\mathcal{F}}$-regular.
4 Explicit representation of conditional Chisini means
In this section we apply Theorem 7 to obtain an explicit formula for conditional Chisini means, which is the cornerstone of the proofs of Theorem 2 and of Corollary 5.
Theorem 8.
Assume that the hypotheses of Theorem 7 are satisfied and consider a resulting pair $(\mathbb{u},\mathbb{P})$ with $\mathbb{u}$ being ${\mathcal{F}}$-regular. Then, for any given a sub-sigma algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\varnothing$ and there exist
  • a ${\mathcal{G}}$-regular function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$, with
@@ 4 Explicit representation of conditional Chisini means
and recall that, under the assumptions of Theorem 8, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is unique up to $\succeq$-null events (see [11], Theorem 3.2 and Proposition 9 in the following section).
To clarify the statement of Theorem 8, it seems important at this stage to point out that the use of $\Phi_{\mathcal{G}}$ in (12) conceals some technical subtleties. To begin with, the dependence of the image of $\mathbb{u}_{\mathcal{G}}$ on $\omega$ requires cautious handling of its domain. Additionally, for $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, the nature of $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ implies that it may not be an element of $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ itself. While the case of a finitely generated $\sigma$-algebra is rather simple (as we shall depict in Example 1 below), we shall dedicate a comprehensive treatment to the issue for an arbitrary $\sigma$-algebra ${\mathcal{G}}$ in the proof postponed to Section 5.3.
Example 1.
@@
For any fixed $A_{i}$ we notice that the image of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$, namely $R_{i}:=\mathrm{Im}(\mathbb{u}_{\mathcal{G}}(\omega,\cdot))$ does not depend on $\omega\in A_{i}$. Given the generalized inverse $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ defined in Theorem 8, it is crucial to observe that the map $x\mapsto\Phi_{\mathcal{G}}(\omega,x)=\mathbb{u}_{\mathcal{G}}^{-1}(\omega,x)$ is well defined as a standard inverse function for $\omega\in A_{i}$ and $x\in R_{i}$ (as a result of strict monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$).
In particular if we now choose any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and set $h:=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ then $h(\omega)\in R_{i}$ for $\omega\in A_{i}$ and indeed
@@
on the set $A_{i}$. This can be readily verified following the procedure in Section 5.3, step 4.
5 Proofs of main results
We stress that the order of the results as presented in the main body of the paper is chosen to put the right focus on the meaning of the results.This section instead is structured in such a way that the logical flow of the proofs is respected.
5.1 Proof of Theorem 4
This section is dedicated to the proof of Theorem 4. We first show that the conditional Chisini mean exists for a preference satisfying (SM), (PC) and (ST) and that it is unique up to $\succeq$-null events. Then we provide a “taking out what is known” property for the conditional Chisini mean before focusing on the proof of the main result.
Proposition 9.
Let $\succeq$ be a preference order satisfying (SM), (PC), (ST) and consider a sub-$\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$ and $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. Then $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\varnothing$ and for any $g,\tilde{g}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ it holds that $\left\{g\neq\tilde{g}\right\}\in{\mathcal{N}}_{\succeq}$.
Proof.
First, observe that for any preference order satisfying (SM) and (PC) there exists a representing functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$. Indeed, (SM) and (PC) impliy the pointwise monotonicity of $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ (see Lemma 18). Consequently, for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ the sets $\underline{F}:=\{x\in\mathbb{R}:f\succeq x\}$ and $\overline{F}:=\{x\in\mathbb{R}:x\succeq f\}$ are non empty closed halflines in $\mathbb{R}$ such that $\overline{F}\cup\underline{F}=\mathbb{R}$. Then $\overline{F}\cap\underline{F}=\left\{z\right\}$, with $z$ being the (unique) certainty equivalent of $f$, i.e. $f\sim z$, and we can define $T(f):=z$.
We show that the hypotheses of [11], Theorem 3.2 hold for such a functional $T$. Indeed, (SM), (PC) and (ST) of $\succeq$ imply that $T$ satisfies [^8] We adopt the notation of [11] (${\mathcal{G}}$-Mo), (${\mathcal{G}}$-PC), (${\mathcal{G}}$-QL) and (${\mathcal{G}}$-NB) for any $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$. It remains to show that $T$ satisfies (${\mathcal{G}}$-PS). Consider $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $A_{1},A_{2}\in{\mathcal{G}}$ with $A_{1}\cap A_{2}=\varnothing$ and let $x_{1},x_{2}\in\mathbb{R}$ be such that $T(f\mathbf{1}_{A_{i}})=T(x_{i}\mathbf{1}_{A_{i}})$ for $i=1,2$. Define $g_{1}:=x_{1}\mathbf{1}_{A_{1}}$ and $g_{2}:=f\mathbf{1}_{A_{2}}$. By (ST) we obtain $T(f\mathbf{1}_{A_{1}}+g_{2}\mathbf{1}_{A_{1}^{c}})=T(x_{1}\mathbf{1}_{A_{1}}+g_{2}\mathbf{1}_{A_{1}^{c}})$ and similarly we find $T(f\mathbf{1}_{A_{2}}+g_{1}\mathbf{1}_{A_{2}^{c}})=T(x_{2}\mathbf{1}_{A_{2}}+g_{1}\mathbf{1}_{A_{2}^{c}})$. Since $A_{2}^{c}\cap A_{1}=A_{1}$ and $A_{1}^{c}\cap A_{2}=A_{2}$, it follows that $T(f\mathbf{1}_{A_{1}}+f\mathbf{1}_{A_{2}})=T(x_{1}\mathbf{1}_{A_{1}}+f\mathbf{1}_{A_{2}})$ and $T(f\mathbf{1}_{A_{2}}+x_{1}\mathbf{1}_{A_{1}})=T(x_{2}\mathbf{1}_{A_{2}}+x_{1}\mathbf{1}_{A_{1}})$, whence $T(f\mathbf{1}_{A_{1}}+f\mathbf{1}_{A_{2}})=T(x_{1}\mathbf{1}_{A_{1}}+x_{2}\mathbf{1}_{A_{2}})$. We can therefore apply [11], Theorem 3.2 which in turn provides $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\varnothing$ and the desired uniqueness. ∎
Lemma 10.
Let $\succeq$ be a preference order satisfying (SM), (PC), (ST) and consider a sub-$\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$ and $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. For $A\in{\mathcal{G}}$, we have
$\mathfrak{m}\left(f\mathbf{1}_{A}\middle|{\mathcal{G}}\right)=\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\mathbf{1}_{A}$ (13)
meaning that $g\in\mathfrak{m}\left(f\mathbf{1}_{A}\middle|{\mathcal{G}}\right)$ if and only if $g=\overline{g}\mathbf{1}_{A}$ for some $\overline{g}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$.
Proof.
Consider $g\in\mathfrak{m}\left(f\mathbf{1}_{A}\middle|{\mathcal{G}}\right)$. By definition we have that $T(g\mathbf{1}_{B})=T((f\mathbf{1}_{A})\mathbf{1}_{B})$ for all $B\in{\mathcal{G}}$. As $\mathbf{1}_{A}\mathbf{1}_{B}=\mathbf{1}_{A\cap B}$ and $A\cap B\in{\mathcal{G}}$, we can write $T((f\mathbf{1}_{A})\mathbf{1}_{B})=T(f\mathbf{1}_{A\cap B})=T(\overline{g}\mathbf{1}_{A\cap B})=T((\overline{g}\mathbf{1}_{A})\mathbf{1}_{B})$, with $\overline{g}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$, and hence it follows that $g=\overline{g}\mathbf{1}_{A}$.
Now let $\overline{g}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ and fix $A\in{\mathcal{G}}$. For all $G\in{\mathcal{G}}$ it holds that $A\cap G\in{\mathcal{G}}$, so that in particular $T(f\mathbf{1}_{A\cap G})=T(\overline{g}\mathbf{1}_{A\cap G})$. It then follows for all $G\in{\mathcal{G}}$ that $T(\overline{g}\mathbf{1}_{A\cap G})=T(f\mathbf{1}_{A\cap G})=T((f\mathbf{1}_{A})\mathbf{1}_{G})=T(g\mathbf{1}_{G}),$ with $g\in\mathfrak{m}\left(f\mathbf{1}_{A}\middle|{\mathcal{G}}\right)$. We conclude that $\overline{g}\mathbf{1}_{A}=g$. ∎
We further observe that, as $\mathfrak{m}\left(0\middle|{\mathcal{G}}\right)=0$ by definition, for any $A\in{\mathcal{G}}$ property (13) is equivalent to
$\mathfrak{m}\left(f\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\middle|{\mathcal{G}}\right)=\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\mathbf{1}_{A}+\mathfrak{m}\left(h\middle|{\mathcal{G}}\right)\mathbf{1}_{A^{c}}.$
Proof of Theorem 4.
We begin by showing that Item (i) implies (ii). To this aim, note that the hypotheses of Proposition 9 are satisfied and hence for any ${\mathcal{G}}\subseteq{\mathcal{F}}$ there exists a function $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ that satisfies Eq. (9).
Item (iii) follows directly from Item (ii).
We conclude the proof showing that (iii) implies (i). Consider $f_{1},f_{2},f,h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and observe that, by definition of ${\mathcal{N}}_{\succeq}$, whenever $A\in{\mathcal{N}}_{\succeq}$ then $f\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\sim h$, so that $f_{1}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\succeq f_{2}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}$ implies $f_{1}\mathbf{1}_{A}+\overline{h}\mathbf{1}_{A^{c}}\succeq f_{2}\mathbf{1}_{A}+\overline{h}\mathbf{1}_{A}^{c}$ for any $\overline{h}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. Similarly whenever $A^{c}\in{\mathcal{N}}_{\succeq}$. Suppose $A,A^{c}\in{\mathcal{F}}\setminus{\mathcal{N}}_{\succeq}$ and consider ${\mathcal{G}}:=\sigma(A)$. First notice that, by construction of ${\mathcal{G}}$ and uniqueness, for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ the conditional Chisini mean $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ collapses to a single ${\mathcal{G}}$-measurable random variable. In particular, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)=x\mathbf{1}_{A}+y\mathbf{1}_{A^{c}}$ for some $x,y\in\mathbb{R}$.
Let $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ be a representing functional for $\succeq$ conditionable on ${\mathcal{G}}=\sigma(A)$ and take $f_{1},f_{2},h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ such that $T(f_{1}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})\geq T(f_{2}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})$. Conditionability of $T$ ensures that there exists $g_{i}\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that $T(g_{i}\mathbf{1}_{G})=T((f_{i}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})\mathbf{1}_{G})$ for any $G\in{\mathcal{G}}$ and $i=1,2$. As we previously pointed out, for this choice of ${\mathcal{G}}$ the conditional Chisini mean consists of a single random variable. Hence we denote with $\mathfrak{m}\left(f_{i}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\middle|{\mathcal{G}}\right)$ the solution $g_{i}$ to the previous equation. By Lemma 10, we have that for $i=1,2$
$\mathfrak{m}\left(f_{i}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\middle|{\mathcal{G}}\right)=\mathfrak{m}\left(f_{i}\middle|{\mathcal{G}}\right)\mathbf{1}_{A}+\mathfrak{m}\left(h\middle|{\mathcal{G}}\right)\mathbf{1}_{A^{c}},$
and recalling that any ${\mathcal{G}}$-measurable function is of the form $x\mathbf{1}_{A}+y\mathbf{1}_{A^{c}}$ for some $x,y\in\mathbb{R}$, we obtain that $\mathfrak{m}\left(f_{i}\middle|{\mathcal{G}}\right)=x_{i}$ on $A$ and analogously $\mathfrak{m}\left(h\middle|{\mathcal{G}}\right)=\overline{y}$ on $A^{c}$ for some $x_{i},\overline{y}\in\mathbb{R}$. Now observe that $T(f_{1}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})\geq T(f_{2}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})$ implies $x_{1}\geq x_{2}$. Assume, by way of contradiction, that $x_{1}<x_{2}$, then (SM) yields $T(x_{1}\mathbf{1}_{A}+\overline{y}\mathbf{1}_{A^{c}})<T(x_{2}\mathbf{1}_{A}+\overline{y}\mathbf{1}_{A^{c}})$. However,
$T(f_{1}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})=T(x_{1}\mathbf{1}_{A}+\overline{y}\mathbf{1}_{A^{c}})<T(x_{2}\mathbf{1}_{A}+\overline{y}\mathbf{1}_{A^{c}})=T(f_{2}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}),$
hence a contradiction.
Keeping $f_{1}$ and $f_{2}$ constant and replacing $h$ with an arbitrary $k\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we find $\mathfrak{m}\left(f_{i}\mathbf{1}_{A}+k\mathbf{1}_{A^{c}}\middle|{\mathcal{G}}\right)=x_{i}\mathbf{1}_{A}+z\mathbf{1}_{A^{c}}$ for $i=1,2$, with $z=\mathfrak{m}(k|{\mathcal{G}})$ on $A^{c}$. As $x_{1}\geq x_{2}$, (SM) implies
$T(f_{1}\mathbf{1}_{A}+k\mathbf{1}_{A^{c}})=T(x_{1}\mathbf{1}_{A}+z\mathbf{1}_{A^{c}})\geq T(x_{2}\mathbf{1}_{A}+z\mathbf{1}_{A^{c}})=T(f_{2}\mathbf{1}_{A}+k\mathbf{1}_{A^{c}}).$
Since $k$ is arbitrary, we get in fact that $T(f_{1}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})\geq T(f_{2}\mathbf{1}_{A}+h\mathbf{1}_{A^{c}})$ for one $h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ entails $T(f_{1}\mathbf{1}_{A}+\overline{h}\mathbf{1}_{A^{c}})\geq T(f_{2}\mathbf{1}_{A}+\overline{h}\mathbf{1}_{A^{c}})$ for all $\overline{h}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, which in turn implies that $\succeq$ satisfies (ST). ∎
5.2 Proof of Theorem 7